User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Katie

Next meeting chair: Earl

Attended: Lim, Prad, Simin, Tiege, Mas, Guifre, Eloise, Jay, Patrick


General Issues

  • Update on Ho checklists? Not yet, but soon.
  • LO Ho 26m problem, looks like it is fixed.
  • Drive controller swapped on the 26m. The problem starting the drives should be fixed now.

Handover Issues

  • r4961 - YG
    • Experiment had to be halted to swap modules. Maybe bad disks?
/home/www/auscope/opswiki/data/pages/operations/agenda2020_09_10.txt · Last modified: 2020/09/10 03:03 by Katie Vandorou